OpenStudy (anonymous):

How do I find the volume of a rectangular box and I do not know what the length is? Just give me the formula. Finding the height is: V = lwh V/lw = lwh/lw V/lw = h Finding the width: ??? Finding the length: ???

OpenStudy (anonymous):

Width and length and the exact same process as height, only you isolate w and l respectively.

OpenStudy (anonymous):

and the length? u isolate w and h?

OpenStudy (jlastino):

w=V/lh l=V/wh

OpenStudy (anonymous):

what? im guessing the length is: V = lwh V/wh = lwh/wh V/wh = l is that it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so the width now is: V = lwh V/lh = lwh/lh V/lh = w correct?

OpenStudy (anonymous):

Yes/

OpenStudy (anonymous):

kthx